### v1

# Large deviations for the maximum of a reducible two-type branching Brownian motion

Abstract.
We consider a two-type reducible branching Brownian motion, defined as a particle system on the real line in which particles of two types move according to independent Brownian motions and create offspring at a constant rate. Particles of type $1$ can give birth to particles of types $1$ and $2$, but particles of type $2$ only give birth to descendants of type $2$. Under some specific conditions, Belloum and Mallein in [4] show that the maximum position $M_{t}$ of all particles alive at time $t$, suitably centred by a deterministic function $m_{t}$, converge weakly. In this work, we are interested in the decay rate of the following upper large deviation probability, as $t\rightarrow\infty$,
$\mathbb{P}(M_{t}\geq\theta m_{t}),\quad\theta>1.$
We shall show that the decay rate function exhibits phase transitions depending on certain relations between $\theta$, the variance of the underling Brownian motion and the branching rate.
Key words and phrases:
Branching Brownian motion, maximum, large deviation, two-type, reducible
2010 Mathematics Subject Classification:
60F10, 60J80
1. Introduction and main results
Branching Brownian motion (BBM) is a spatial branching process of great interests in recent years. On the one hand, its connection with the partial differential Fisher-Kolmogorov-Petrovsky-Piskunov (F-KPP) equation brings out many studies on both probabilistic and analytic sides, see for instance, [1, 18, 10, 16, 3]. On the other hand, it is the fundamental model to understand the BBM-universality class which includes the 2-dim Gaussian free field [8, 9] and 2-dim cover times [12], etc.
An one-dimensional standard binary branching Brownian motion is a continuous-time particle system on the real line which can be constructed as follows. It starts with one individual located at the origin at time 00 that moves according to a Brownian motion with variance $\sigma^{2}>0$. After an independent exponential time of parameter $\beta$, the initial particle dies and gives birth to 2 children that start at the position their parent occupied at its death. These $2$ children then move according to independent Brownian motions with variance $\sigma^{2}$ and give birth independently to their own children at rate $\beta$. The particle system keeps evolving in this fashion for all time.
In this work, we shall study the two-type reducible branching Brownian motion. This is a particle system on the real line in which particles possess a type in addition with their position. Particles of type $1$ move according to Brownian motions with variance $\sigma^{2}$ and branch at rate $\beta$ into two children of type $1$. Additionally, they give birth to particles of type $2$ at rate $\alpha$. Particles of type $2$ move according to Brownian motions with variance $\sigma_{2}^{2}$ and branch at rate $\beta_{2}$, but cannot give birth to descendants of type $1$. Up to a dilation of time and space, in this work we shall always assume that $\beta_{2}=\sigma_{2}=1$. This model have been studied extensively by many authors; see [7, 14, 15, 17].
With abuse of notation, for all $t\geq 0$, we denote by ${\mathcal{N}}_{t}$ the collection of the individuals alive at time $t$ for both the standard BBM and the two-type reducible BBM. And denote by ${\mathcal{N}}_{t}^{1}$ and ${\mathcal{N}}_{t}^{2}$ the set of particles of type 1 and type 2 at time $t$, respectively. For any $u\in{\mathcal{N}}(t)$ and $s\leq t$, let $X^{u}_{s}$ denote the position at time $s$ of the individual $u$ or its ancestor alive at that time. The maximum of the branching Brownian motion at time $t$ is defined as
$M_{t}:=\max\{X^{u}_{t}:u\in{\mathcal{N}}(t)\}.$
To distinguish the two-type reducible BBM and the standard BBM, we shall regard the standard BBM as $(\sigma^{2},\beta)$-BBM to emphasis the variance of the Brownian motion and the branching rate. We also denote by $\mathbb{P}_{\sigma^{2},\beta}$ its law and write $\bar{X}^{u}_{s}$ for the position at time $s$ of the individual $u$ of a $(\sigma^{2},\beta)$-BBM.
For $(\sigma^{2},\beta)$-BBM, the asymptotic behaviour of
$\bar{M}_{t}:=\max\{\bar{X}_{t}^{u}:u\in{\mathcal{N}}(t)\},$
as $t\to\infty$, has been subjected to intense study, partly due to its link with the F-KPP reaction-diffusion equation, defined as
(1) $\partial_{t}u(t,x)=\frac{1}{2}\partial^{2}_{x}u(t,x)+(1-u(t,x))-(1-u(t,x))^{2},\quad t\geq 0,\,x\in{\mathbb{R}}.$
Precisely, McKean [18] showed that
$\mathbb{P}_{\sigma^{2},1}(\bar{M}_{t}\geq\sigma x)=\bar{u}(t,x),$
where the function $(t,x)\mapsto\bar{u}(t,x)$ is the unique solution of (1) with initial condition $\bar{u}(0,x)=\mathbf{1}_{\{x<0\}}$. With the help of (1), Bramson [10] proved that for a $(\sigma^{2},\beta)$-BBM,
$\bar{M}_{t}-\sqrt{2\sigma^{2}\beta}t+\frac{3}{2\sqrt{2\beta/\sigma^{2}}}\ln t,$
converges weakly, as $t\rightarrow\infty$. In particular, uniformly in $x\in\mathbb{R}$, as $t\rightarrow\infty$,
(2) $1-\bar{u}(t,\bar{m}_{t}+x)=\mathbb{P}_{\sigma^{2},1}\left(\frac{\bar{M}_{t}}{\sigma}-\sqrt{2}t+\frac{3}{2\sqrt{2}}\ln t\leq x\right)\longrightarrow\bar{\omega}(x),$
where $\bar{\omega}$ is the travelling wave solution of the F-KPP equation at speed $\sqrt{2}$, which satisfies
$\frac{1}{2}\bar{\omega}^{\prime\prime}+\sqrt{2}\bar{\omega}^{\prime}-\bar{\omega}(1-\bar{\omega})=0.$
Lalley and Sellke [16] characterised $\bar{\omega}$ in terms of a random shift of the Gumbel distribution by introducing the so-called derivative martingale.
Parallel results of (2) for the two-type reducible BBM were obtained in [4]. To state their results, we need to decompose the state space $(\beta,\sigma^{2})\in{\mathbb{R}}_{+}^{2}$ into three regions:
$\mathcal{C}_{\text{I}}$ =$=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}>\frac{\mathbb{1}_{\{\beta\leq 1\}}}{\beta}+\mathbb{1}_{\{\beta>1\}}\frac{\beta}{2\beta-1}\right\},$
$\mathcal{C}_{\text{II}}$ =$=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}<\frac{\mathbb{1}_{\{\beta\leq 1\}}}{\beta}+\mathbb{1}_{\{\beta>1\}}(2-\beta)\right\},$
$\mathcal{C}_{\text{III}}$ =$=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}+\beta>2\text{ and }\sigma^{2}<\frac{\beta}{2\beta-1}\right\}.$
From Belloum and Mallein [4], we have:
$M_{t}-m_{t}$
converges weakly, as $t\rightarrow\infty$, where
$m_{t}=\begin{cases}\sqrt{2\sigma^{2}\beta}t-\frac{3}{2\sqrt{2\beta/\sigma^{2}}}\ln t,\quad&(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}},\cr\sqrt{2}t-\frac{3}{2\sqrt{2}}\ln t,\quad&(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}},\cr\frac{\sigma^{2}-\beta}{\sqrt{2\left(1-\sigma^{2}\right)(\beta-1)}}t,\quad&(\beta,\sigma^{2})\in\mathcal{C}_{\text{III}}.\end{cases}$
From above, we see that if $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{I}$, the speed of the two-type reducible BBM is $\sqrt{2\beta\sigma^{2}}$, which is the same as particle s of type $1$ alone, ignoring births of particles of type $2$. Conversely, if $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$, then the speed of the process is $\sqrt{2}$, equal to the speed of a single BBM of particles of type $2$. Finally, if $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$, the speed of the process is a mixture of the long-time asymptotic of the processes of particles of type $1$ and $2$, which is larger than $\max\left(\sqrt{2},\sqrt{2\beta\sigma^{2}}\right)$.
In this article, we study the large deviation probabilities for the maximum of the two-type reducible BBM. That is the convergence rate of the following probability
$\mathbb{P}(M_{t}>\theta m_{t}),\quad\theta>1,$
as $t\rightarrow\infty$. We divide the main result of our article into three theorems, one for each area the pair $\left(\beta,\sigma^{2}\right)$ belongs to.
Theorem 1.1.
If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$, then the following limit exists for any $\theta>1$,
$\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}}\theta t\right)=A,$
where for $\beta<1$,
$A=\beta\left(1-\theta^{2}\right),$
for $\beta>1$ and $\sigma^{2}\geq 1$, we have
$A=\beta\left(1-\theta^{2}\right),$
for $\beta>1$ and $\sigma^{2}<1$, we have
$A=\begin{cases}\beta\left(1-\theta^{2}\right),&\theta^{2}\leq\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta},\\ 1-\theta^{2}\beta\sigma^{2},&\theta^{2}\geq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta},\\ \frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.\end{cases}$
Theorem 1.2.
If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$, then the following limits exists for any $\theta>1$,
$\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)=:A,$
where for $\beta>1$, we have
$A=1-\theta^{2},$
for $\beta\leq 1,\sigma^{2}\leq 1$, we have
$A=1-\theta^{2},$
for $\beta\leq 1,\sigma^{2}>1$, we have
$A=\begin{cases}1-\theta^{2},&\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)},\cr\beta-\frac{\theta^{2}}{\sigma^{2},}&\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4},\cr\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}<\theta^{2}<\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.\end{cases}$
Theorem 1.3.
If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$, then for any $\theta>1$, we have
$\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\xi\theta t\right)=\begin{cases}1-\frac{\xi^{2}\theta^{2}}{2},&\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}},\\ \beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}},&\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}},\cr\frac{\sqrt{2}(1-\beta)-\theta^{2}\left(\sigma^{2}-\beta\right)^{2}-1}{\sqrt{2}\left(\sigma^{2}-1\right)}+1,&\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}},\end{cases}$
where $\xi=\frac{\sigma^{2}-\beta}{\sqrt{2(1-\sigma^{2})(\beta-1)}}.$
Let us mention here that in the context of standard $(1,1)$-BBM (i.e., $\sigma^{2}=1,\beta=1$), one would have: For any $\theta>1$ and $x\in{\mathbb{R}}$,
(3) $\lim_{t\rightarrow\infty}\sqrt{t}e^{(\theta^{2}-1)t}\mathbb{P}_{1,1}(\bar{M}_{t}\geq\sqrt{2}\theta t+x)$ $=\lim_{t\rightarrow\infty}\sqrt{t}e^{(\theta^{2}-1)t}\bar{u}(t,\sqrt{2}\theta t+x)$
(4) $=e^{\sqrt{2}x}C(\theta,\bar{u}),$
where $C(\theta,\bar{u})$ is a constant depending only on $\theta$ and $\bar{u}$; see [11], [2] and [5]. In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of two-type inreduciable BBM. It is possible to obtain more preciser form as (3) with the help of truncated second moment method and the F-KPP equation. Our method may also apply to study the cases of $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{II}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}}\cap\mathcal{C}_{\text{III}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{III}}$; see [17] for related results.
In the next section, we will present some technical lemmas borrowed from existed literatures. The proofs for Theorems 1.1, 1.2 and 1.3 will be given in Sections 3, 4 and 5, respectively.
2. Technical lemmas
Lemma 2.1.
([13], Lemma 2.1) Let $X$ be a centered Gaussian random variables with variance $\sigma^{2}>0$. Then
(5) $\mathbb{P}(X>a)=(2\pi)^{-1/2}(\sigma/a)e^{-\frac{a^{2}}{2\sigma^{2}}}\left(1+O(\sigma^{2}/a^{2})\right),\quad\text{as }a/\sigma\rightarrow\infty,$
with the r.h.s. above without error term being an upper bound valid for any $a>0.$
Furthermore, for $(\sigma^{2},\beta)$-BBM and $10<x<$ 1, $\mathbb{P}_{\sigma^{2},\beta}-a.s.$,
(6) $\frac{1}{t}\log{\mathcal{N}}_{t}\left(\sqrt{2\beta}x\sigma t)\right)\overset{t\rightarrow\infty}{\longrightarrow}(1-x^{2})\beta,$
where ${\mathcal{N}}_{t}(x)$ denote the number of particles, in the $(\sigma^{2},\beta)$-BBM, alive at time $t$ and positioned in $[tx,\infty)$; see [6].
If $u\in\mathcal{N}_{t}^{2}$, we denote by $T(u)$ the time at which the oldest ancestor of type 2 of $u$ was born. The set $\mathfrak{B}$ of particles of type $2$ that are born from a particle of type $1$, that can be defined as
$\mathfrak{B}:=\{u\in\cup_{t\geq 0}{\mathcal{N}}^{2}_{t}:T(u)=b_{u}\},$
where $b_{u}$ is the birth time of $u$. We observe that $\mathfrak{B}$ can be thought of as a Poisson point process with random intensity.
Lemma 2.2.
([4], Lemma 4.2) Set ${\mathcal{F}}_{1}=\sigma(X_{t}^{u},u\in{\mathcal{N}}_{t}^{1},t\geq 0)$. The point measure $\sum_{u\in\mathfrak{B}}\delta_{(X_{s}^{u},s\leq T(u))}$ conditionally on ${\mathcal{F}}_{1}$ is a Poisson point process with intensity $\alpha{\rm d}t\times\sum_{u\in{\mathcal{N}}_{t}^{1}}\delta_{(X_{s}^{u},s\leq t)}$.
The results in the following proposition are variations of the so-called many-to-one lemma.
Lemma 2.3.
For all measurable non-negative functionals $f$ and $g$, we have
$\mathbb{E}\left[\sum_{u\in{\mathcal{N}}_{t}^{1}}f(X_{s}^{u},s\leq t)\right]=e^{\beta t}\mathbb{E}\left[f(\sigma B_{s},s\leq t)\right]$
and
$\mathbb{E}\left[\sum_{u\in{\mathcal{N}}_{t}^{2}}g((X_{s}^{u},s\leq t),T(u))\right]$
$\quad=\alpha\int_{0}^{t}e^{\beta s+(t-s)}E\left[g((\sigma B_{u\wedge s}+(B_{u}-B_{u\wedge s}),u\leq t),s)\right]{\rm d}s.$
Proof.
The first result is just the usual many-to-one lemma. The second one can be found in Proposition 4.1 in [4]. ∎
3. Proof of Theorem 1.1
In this section, we consider the case that
$(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}>\frac{\mathbb{1}_{\{\beta\leq 1\}}}{\beta}+\frac{\beta\mathbb{1}_{\{\beta>1\}}}{2\beta-1}\right\}.$
3.1. Upper bound
Proof.
From Lemma 2.3, we have
(7) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
(8) $=$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
(9) $\leqslant$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)+\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
(10) $\leqslant$ $2\max\left[e^{\beta t}\mathbb{P}\left(\sigma B_{t}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right),\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{P}\left(\sigma B_{s}+B_{t}-B_{s}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)\mathrm{d}s\right]$
(11) =:absent:=:= : $2\max\{I_{1},I_{2}\}.$
Then by (5)
$I_{1}$ $\leq O(1)\cdot\frac{1}{\sqrt{t}}\cdot e^{\beta t}\cdot e^{-\frac{2\beta\theta^{2}t^{2}}{2t}}$
$=\frac{O(1)}{\sqrt{t}}\cdot e^{\beta\left(1-\theta^{2}\right)t}.$
Noticing that
$\operatorname{Var}\left(\sigma B_{s}+B_{t}-B_{s}\right)=\sigma^{2}\cdot s+(t-s),$
we have
(12) $I_{2}$ $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\sigma^{2}\beta\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}\mathrm{d}s$
(13) $=\alpha\int_{0}^{1}e^{f(u)t}\cdot\frac{\sqrt{t}}{\sqrt{2\pi\left((\sigma^{2}-1)u+1\right)}}\mathrm{d}u,$
where
(14) $f_{1}(u)=(\beta-1)u+1-\frac{\sigma^{2}\beta\theta^{2}}{(\left.\sigma^{2}-1\right)u+1}.$
Notice that $f_{1}(1)=\beta(1-\theta^{2})$. Then our task is reduced to know $\max_{0\leq u\leq 1}f(u)$. To end this, let us consider
$f_{1}^{\prime}(u)=\beta-1+\frac{\sigma^{2}\beta\theta^{2}}{\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}\left(\sigma^{2}-1\right).$
We have three cases.
i). If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\leq 1$, then $u\mapsto f_{1}^{\prime}(u)$ is an increasing function. In this case, by noticing that $\beta\sigma^{2}>1$, one has
$f_{1}^{\prime}(0)=\beta-1+\sigma^{2}\beta(\sigma^{2}-1)=\beta(1-\sigma^{2})+\sigma^{2}\beta\sigma^{2}-1>\beta(1-\sigma^{2})+\sigma^{2}-1\geq 0\cdot$
Thus $f_{1}^{\prime}(u)\geq 0$ for all $u\in[0,1]$. Hence $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and then we have
(15) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
ii). If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\geq 1,\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
(16) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
iii). If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
First, if $\theta^{2}<\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}$, which implies $f^{\prime}(1)>0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and
(17) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
Second, if $\theta^{2}>\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$,  which means $f_{1}^{\prime}(0)<0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(0)$ and
(18) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq 1-\theta^{2}\beta\sigma^{2}.$
Last, if $\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$, then $f_{1}^{\prime}(0)\geq 0$ and $f^{\prime}(1)\leq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$, where $f_{1}^{\prime}(u_{0})=0$. And one can check that
(19) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\sigma\theta\cdot\sqrt{\beta}-1}{\left|\sigma^{2}-1\right|},$
and
$f_{1}(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
Therefore,
(20) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
This proves the upper bound. ∎
3.2. Lower bound for Theorem 1.1
Notice that
$\mathbb{P}\left(M_{t}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)\right\}.$
First, from 3, we have
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$=\beta\left(1-\theta^{2}\right).$
On the other hand, set
$\mathfrak{B}_{t}=\{u\in\mathfrak{B}:T(u)<t\}.$
According to construction of our model, we have
(21) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(22) $=\mathbb{P}\left(\exists u\,\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(23) $=1-\mathbb{P}\left(\forall\,u\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}<\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(24) $=1-\mathbb{E}\left[\prod_{u\in\mathfrak{B}_{t}}\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right]$
(25) $=1-\mathbb{E}\left[\exp\left(-\sum_{u\in\mathfrak{B}_{t}}-\ln\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right)\right],$
where
$1-\varphi_{t}(x,s)=\bar{u}\left(t-s,\,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x\right)=\mathbb{P}_{1,1}\left(\bar{M}_{t-s}\geq\sqrt{2\sigma^{2}\theta^{2}\beta}\cdot t-x\right).$
and $\left\{M^{u}_{t-T(u)}:u\in\mathfrak{B}_{t}\right\}$ are independent random variables such that $M^{u}_{t-s}$ has the same law as $\bar{M}_{t-s}$ under $\mathbb{P}_{1,1}$. Applying lemma 2.2, we have
(26) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(27) $=1-\mathbb{E}\left[\exp\left\{-\alpha\int_{0}^{t}\sum_{u\in{\mathcal{N}}_{s}^{1}}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right\}\right]$
(28) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right\}\right]$
(29) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right]$
(30) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right],$
where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$, and
${\mathcal{N}}_{s}^{1}(a)=\{u\in{\mathcal{N}}_{s}^{1}:X_{s}^{u}\geq a\}.$
Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Furthermore, by Jensen’s inequality and Fubini’s theorem, we have
(31) $\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right]$
(32) $\leq\int_{I_{t}}\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s.$
3.2.1. The case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(0)$.
Choose
$I_{t}=[\varepsilon t,\varepsilon t+1],\quad x_{t}=-\varepsilon t,$
for $\varepsilon>0$ small enough. Then for $s\in I_{t}$,
$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(-\varepsilon t)}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\},|{\mathcal{N}}_{s}^{1}(-\varepsilon t)|\geq 1\right]$
$\quad+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(-\varepsilon t)|=0\right)$
$\leq\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right)+\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}<-\varepsilon t\right).$
Therefore,
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$\geq 1-\int_{I_{t}}\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s$
$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right){\rm d}s$
$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-\varepsilon t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right){\rm d}s.$
Then applying (3), we have
$\lim_{\varepsilon\rightarrow 0}\lim_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\geq 1-\sigma^{2}\beta\theta^{2},$
uniformly for $s\in[0,1]$. We further notice that
$\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right)\rightarrow 1,\quad\text{as }t\rightarrow\infty,$
uniformly for $s\in[0,1]$. Then applying Jensen’s inequality and Fatou’s lemma, one has
$\liminf_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$ $=1-\sigma^{2}\beta\theta^{2}$
$\geq\beta(1-\theta^{2}),$
where the last inequality holds if $\theta^{2}\geq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\beta}$. This proves Theorem 1.1 for the case of $\theta^{2}\geq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$, when $\sigma^{2}<1$ and $\beta>1$.
3.2.2. The case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$.
This case corresponds to the case that
$\left\{\left(\theta^{2},\,\beta,\,\sigma^{2}\right):\theta^{2}>1,\,\beta\leq 1,\,\sigma^{2}>\frac{1}{\beta}\right\}\bigcup\left\{\left(\theta^{2},\,\beta,\,\sigma^{2}\right):\theta^{2}>1,\,\beta>1,\,\sigma^{2}>1\right\}$
$\quad\bigcup\left\{\left(\theta^{2},\beta,\sigma^{2}\right):\theta^{2}\leq\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta},\,\beta>1,\,\frac{\beta}{2\beta-1}<\sigma^{2}<1\right\}.$
Choose
$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2\beta\sigma^{2}}\theta t+\varepsilon t,$
for $\varepsilon>0$ small enough. Then for $s\in I_{t}$,
$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,-\varepsilon t\right)\right\},|{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right]$
$\quad+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|=0\right).$
Notice that
$\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)=\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right).$
Therefore,
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right){\rm d}s$
$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\bar{u}\left(t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right){\rm d}s$
$=\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>x_{t}\right){\rm d}s.$
Since
$\lim_{t\rightarrow\infty}\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)=1$
and
$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>x_{t}\right)=\beta(1-\theta^{2}),$
uniformly for $s\in[0,1]$, then we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)=\beta(1-\theta^{2}).$
3.2.3. The case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$.
Notice that where $f_{1}^{\prime}(u_{0})=0$. This corresponds to the case that $\beta>1,\sigma^{2}<1$, and
$\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.$
We also recall that
$f(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
For $\varepsilon>0$ small enough, choose
$I_{t}=[u_{0}t,u_{0}t+1],\quad x_{t}=x_{0}t,$
for some $0<x_{0}<\sqrt{2\beta\theta^{2}\sigma^{2}}$, whose value will be determined later. Then for $s\in I_{t}$,
$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\},|{\mathcal{N}}_{s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right]$
$\quad+\mathbb{P}_{\sigma^{2},\beta}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|<e^{(\beta_{0}-\varepsilon)t}\right)$
$\leq\exp\left\{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\}\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)$
$\quad+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|<e^{(\beta_{0}-\varepsilon)t}\right),$
where
$\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0,\quad\text{and}\quad x_{0}=\frac{\sqrt{2\beta}\theta u_{0}\sigma^{3}}{1-u_{0}+u_{0}\sigma^{2}}.$
Therefore,
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\}\right)$
×$\qquad\quad\times\mathbb{P}_{\sigma^{2},\beta}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
We also notice that
$\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}>1,$
since
$\sigma^{2}\geq\frac{\beta}{2\beta-1}>2-\beta,$
for $\beta>1$. Then applying (3) again, we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)$
$\quad\geq\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
uniform for $s\in[0,1]$. Furthermore, according to (6),
$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
Thus
$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1.$
Since $\varepsilon>0$ is arbitrary, then applying Jensen’s inequality and Fatou’s lemma, one has
(33) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
(34) $\geq\beta_{0}+\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
(35) $=(\beta-1)u_{0}+1-\frac{\sigma^{2}\beta\theta^{2}}{(\left.\sigma^{2}-1\right)u_{0}+1}$
(36) $=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}$
(37) $\geq\beta(1-\theta^{2})=f_{1}(1),$
where the last inequality holds for the case that $\beta>1,\sigma^{2}<1$, and
$\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.$
This proves Theorem 1.1 for the above case.
4. Proof of Theorem 1.2
In this section, we always suppose that
$(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}}=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}<\frac{\mathbb{1}_{\{\beta\leq 1\}}}{\beta}+\mathbb{1}_{\{\beta>1\}}(2-\beta)\right\}.$
The arguments are exactly the same to those in Section 3. One just to replace $\sqrt{2\beta\sigma^{2}}\theta$ by $\sqrt{2}\theta$. We only give an outline here.
4.1. Upper bound for Theorem 1.2
We first consider the upper bound.
Proof.
First,
$\mathbb{P}\left(M_{t}\geqslant\sqrt{2}\theta t\right)$
[…]
$=:2\max(\text{I},\,\text{II}).$
By (5), we have
$\text{I}\leq O(1)\cdot\frac{1}{\sqrt{t}}\cdot e^{\beta t}\cdot e^{-\frac{2\theta^{2}t^{2}}{2\sigma^{2}t}}=\frac{O(1)}{\sqrt{t}}\cdot e^{\left(\beta-\frac{\theta^{2}}{\sigma^{2}}\right)t}.$
and
$\text{II}\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}.$
Set
(38) $f_{2}(u)=(\beta-1)u+1-\frac{\theta^{2}}{(\left.\sigma^{2}-1\right)u+1},\quad u\in(0,1).$
Then
$f_{2}^{\prime}(u)=\beta-1+\frac{\theta^{2}}{\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}\left(\sigma^{2}-1\right).$
Solving $f^{\prime}(u_{0})=0$, we have
(39) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\theta-1}{\left|\sigma^{2}-1\right|},$
and
$f_{2}(u_{0})=\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
Then we have several subcases.
Case I. $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$ and $\beta>1$.
In this case $u\mapsto f_{2}^{\prime}(u)$ is increasing and $f_{2}^{\prime}(1)<0$. Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(0)$ and hence
(40) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
Case II. $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$ and $\beta\leq 1,\sigma^{2}<1$.
In this case $f_{2}^{\prime}(u)<0$ for all $u\in[0,1]$, Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(0)$ and hence
(41) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
Case III. $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$ and $\beta\leq 1,\sigma^{2}\geq 1$. In this case, $u\mapsto f_{2}^{\prime}(u)$ is decreasing.
When $\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}$, then $f_{2}^{\prime}(0)<0$. Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(0)$ and hence
(42) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
When $\theta^{2}>\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}$, then $f_{2}^{\prime}(1)>0$, Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(1)$ and hence
(43) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq\beta-\frac{\theta^{2}}{\sigma^{2}}.$
When $\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}\leq\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}$, then $f_{2}^{\prime}(0)\geq 0$ and $f_{2}^{\prime}(1)\leq 0$. Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(u_{0})$ and hence
(44) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
4.2. Lower bound for Theorem 1.2
Notice that
$\mathbb{P}\left(M_{t}\geqslant\sqrt{2}\theta t\right)$
$\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\right\}.$
First, from 3, we have
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2}\theta t\right)$
$=\beta-\frac{\theta^{2}}{\sigma^{2}}.$
By similar reasoning as (21), (26) and (31), we have
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
$\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s,$
where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$. Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Recall from (38) that
$f_{2}(u)=(\beta-1)u+1-\frac{\theta^{2}}{(\left.\sigma^{2}-1\right)u+1},\quad u\in(0,1).$
4.2.1. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(0)$.
This corresponds to the cases that $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, or $\beta\leq 1,\sigma^{2}>1$ and $\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}.$ The arguments are parallel to those in Section 3.2.1. We only give an outline. Choose
$I_{t}=[\varepsilon t,\varepsilon t+1],\quad x_{t}=-\varepsilon t,$
for $\varepsilon>0$ small enough. Then
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-\varepsilon t-s,\sqrt{2}\theta t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right){\rm d}s.$
Then applying (3), we have
$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-s,\sqrt{2}\theta t+\varepsilon t\right)\geq 1-\theta^{2}.$
uniformly for $s\in[0,1]$. We further notice that
$\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right)\rightarrow 1,\quad\text{as }t\rightarrow\infty,$
uniformly for $s\in[0,1]$. Then applying Jensen’s inequality and Fatou’s lemma, one has
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $\geq 1-\theta^{2}.$
This proves Theorem 1.2 for the cases that $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, or $\beta\leq 1,\sigma^{2}>1$ and $\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}$, since $1-\theta^{2}\geq\beta-\frac{\theta^{2}}{\sigma^{2}}$ in all above cases.
4.2.2. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(1)$.
This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Again, we only give an outline of the proof. Choose
$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2}\theta t+\varepsilon t,$
for $\varepsilon>0$ small enough. Then we have
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\sqrt{2}\theta t+\varepsilon t\right){\rm d}s.$
Since
$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(M_{(1-\varepsilon)t+s}>\sqrt{2}\theta t+\varepsilon t\right)\geq\beta-\frac{\theta^{2}}{\sigma^{2}},$
then we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\geq\beta-\frac{\theta^{2}}{\sigma^{2}}.$
This proves Theorem 1.2 for the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$
4.2.3. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(u_{0})$ for $10<u_{0}<$ 1.
Notice that $f_{2}^{\prime}(u_{0})=0$. This corresponds to the case that $\beta<1,1<\sigma^{2}<\frac{1}{\beta}$, and
$\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}<\theta^{2}<\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$
We also recall that in this case
$10<u_{0}=\frac{\sqrt{\frac{\sigma^{2}-1}{1-\beta}}\cdot\theta-1}{\sigma^{2}-1}<$ 1
and
$f_{2}(u_{0})=\frac{-2\sqrt{(1-\beta)\left(\sigma^{2}-1\right)}\theta-\beta+\sigma^{2}}{\sigma^{2}-1}.$
Choose
$x_{0}=\frac{\sqrt{2}\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t,\quad\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0.$
One can check that $x_{0}<\sqrt{2}\theta$ since $u_{0}>0.$ Then,
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
$\geq\int_{0}^{1}\left(1-e^{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\sqrt{2}\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
We also notice that
$\frac{\left(\sqrt{2}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}>1,$
since
$\sigma^{2}<\frac{1}{\beta}<2-\beta,\quad\text{for}\quad\beta<1.$
Then applying (3) again, we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\sqrt{2}\theta t-x_{0}t\right)\geq\left(1-\frac{\left(\sqrt{2}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
uniform for $s\in[0,1]$. Furthermore, according to (6),
$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
Thus
$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1,$
uniform for $s\in(0,1)$. Then one has
(45) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
(46) $\geq\beta_{0}+\left(1-\frac{\left(\sqrt{2}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
(47) $=(\beta-1)u_{0}+1-\frac{\theta^{2}}{\left(\sigma^{2}-1\right)u_{0}+1}$
(48) $=\frac{-2\sqrt{(1-\beta)\left(\sigma^{2}-1\right)}\theta-\beta+\sigma^{2}}{\sigma^{2}-1}$
(49) $\geq\beta-\frac{\theta^{2}}{\sigma^{2}}=f_{2}(1),$
where the last inequality holds for the case that $\beta<1,\sigma^{2}>1$, and
$\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\leq\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$
This proves Theorem 1.2 for the above case.
5. Proof of Theorem 1.3
In this section, we always assume that
$(\beta,\sigma^{2})\in\mathcal{C}_{\text{III}}=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}+\beta>2\text{ and }\sigma^{2}<\frac{\beta}{2\beta-1}\right\}.$
Recall $\xi=\frac{\sigma^{2}-\beta}{\sqrt{2(1-\sigma^{2})(\beta-1)}}.$ The arguments are also exactly the same to those in Section 4. One just to replace $\theta$ by $\frac{\xi\theta}{\sqrt{2}}$. Again, We only give an outline here.
5.1. Upper bound
Proof.
First, by union bound, one has
$\mathbb{P}\left(M_{t}\geqslant\xi\theta t\right)$
$\leqslant$ $2\max\left(e^{\beta t}\mathbb{P}\left(\sigma B_{t}\geqslant\xi\theta t\right),\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{P}\left(\sigma B_{s}+B_{t}-B_{s}\geqslant\xi\theta t\right){\rm d}s\right)$
=:absent:=:= : $2\max\left(\text{I},\,\text{II}\right).$
According to (5), we have
$\text{I}\leq O(1)\cdot\frac{1}{\sqrt{t}}\cdot e^{\beta t}\cdot e^{-\frac{\xi^{2}\theta^{2}t^{2}}{2\sigma^{2}t}}=\frac{O(1)}{\sqrt{t}}\cdot e^{\left(\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}\right)t}$
and
$\text{II}\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{\xi^{2}\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}.\$
For II, set
(50) $f_{3}(u)=(\beta-1)u+1-\frac{\xi^{2}\theta^{2}}{2((\left.\sigma^{2}-1\right)u+1)},\quad u\in(0,1).$
Then
$f_{3}^{\prime}(u)=\beta-1+\frac{\xi^{2}\left(\sigma^{2}-1\right)\theta^{2}}{2\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}.$
Solving $f^{\prime}(u_{0})=0$ yields
(51) $u_{0}=\frac{\xi\theta\sqrt{\frac{1-\sigma^{2}}{2(\beta-1)}}-1}{\sigma^{2}-1}$
and
$f_{3}(u_{0})=\frac{\sqrt{2}(1-\beta)-\theta^{2}\left(\sigma^{2}-\beta\right)^{2}-1}{\sqrt{2}\left(\sigma^{2}-1\right)}+1.$
Since $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$, then $\beta>1,\sigma^{2}<1$ and hence $u\mapsto f_{3}^{\prime}(u)$ is decreasing. We have three cases.
If $\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, then $f_{3}^{\prime}(0)\leq 0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(0)$ and hence
(52) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\xi\theta t\right)\leq f_{3}(0)=1-\frac{\xi^{2}\theta^{2}}{2}.$
If $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$, then $f^{\prime}(1)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(1)$ and hence
(53) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\xi\theta t\right)\leq f_{3}(1)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
If $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, then $f_{3}^{\prime}(0)>0$ and $f_{3}^{\prime}(1)<0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ and hence
(54) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln P\left(M_{t}>\xi\theta t\right)\leq f_{3}(u_{0})=\frac{\sqrt{2}(1-\beta)-\theta^{2}\left(\sigma^{2}-\beta\right)^{2}-1}{\sqrt{2}\left(\sigma^{2}-1\right)}+1.$
5.2. Lower bound for Theorem 1.3
Notice that
$\mathbb{P}\left(M_{t}\geqslant\xi\theta t\right)\geqslant\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right),\,\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)\right\}.$
First, from 3, we have
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\xi\theta t\right)$
$=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
By similar reasoning as (21), (26) and (31), we have
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$
$\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\xi\theta t-x_{t}\right)\right\}\right]{\rm d}s,$
where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$. Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Recall from (50) that
$f_{3}(u)=(\beta-1)u+1-\frac{\xi^{2}\theta^{2}}{2((\left.\sigma^{2}-1\right)u+1)},\quad u\in(0,1).$
5.2.1. The case of $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(0)$.
This corresponds to the cases that $\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$. Choose
$I_{t}=[\varepsilon t,\varepsilon t+1],\quad x_{t}=-\varepsilon t,$
for $\varepsilon>0$ small enough. Then
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$
$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-\varepsilon t-s,\xi\theta t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right){\rm d}s.$
Then applying (3), we have
$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-s,\xi\theta t+\varepsilon t\right)\geq 1-\frac{\xi^{2}\theta^{2}}{2}.$
We further notice that
$\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right)\rightarrow 1,\quad\text{as }t\rightarrow\infty,$
Then applying Jensen’s inequality and Fatou’s lemma, one has
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $\geq 1-\frac{\xi^{2}\theta^{2}}{2}.$
This proves Theorem 1.3 for the cases that $\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, since $1-\frac{\xi^{2}\theta^{2}}{2}\geq\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}$ in this case.
5.2.2. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(1)$.
This case corresponds to the case that $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$. Choose
$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2}\theta t+\varepsilon t,$
for $\varepsilon>0$ small enough. Then we have
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$
$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\xi\theta t+\varepsilon t\right){\rm d}s.$
Since
$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\xi\theta t+\varepsilon t\right)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}},$
then we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
This proves Theorem 1.3 for the case that $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$.
5.2.3. The case of $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ for $10<u_{0}<$ 1.
Notice that $f_{3}^{\prime}(u_{0})=0$. This corresponds to the case that $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$. Choose
$x_{0}=\frac{\xi\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t,\quad\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0.$
Obviously $x_{0}<\xi\theta$ since $u_{0}>0.$ Then
$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$
$\geq\int_{0}^{1}\left(1-e^{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\,\xi\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
We also notice that
$\frac{\left(\xi\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}>1.$
Then applying (3) again, we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\xi\theta t-x_{0}t\right)\geq\left(1-\frac{\left(\xi\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
uniform for $s\in[0,1]$. Furthermore, according to (6),
$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
Thus
$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1,$
uniform for $s\in(0,1)$. Then one has
(55) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=f_{3}(u_{0})$
(56) $\geq\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}=f_{3}(1),$
where the last inequality holds for the case that $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$. This proves Theorem 1.3 for the above case. We are done.

### v2

@@
 Particles of type $1$ can give birth to particles of types $1$ and $2$, but particles of type $2$ only give birth to descendants of type $2$.
-Under some specific conditions, Belloum and Mallein in [4] show that the maximum position $M_{t}$ of all particles alive at time $t$, suitably centred by a deterministic function $m_{t}$, converge weakly.
+Under some specific conditions, Belloum and Mallein in [4] showed that the maximum position $M_{t}$ of all particles alive at time $t$, suitably centered by a deterministic function $m_{t}$, converge weakly.
 In this work, we are interested in the decay rate of the following upper large deviation probability, as $t\rightarrow\infty$,
@@
 Finally, if $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$, the speed of the process is a mixture of the long-time asymptotic of the processes of particles of type $1$ and $2$, which is larger than $\max\left(\sqrt{2},\sqrt{2\beta\sigma^{2}}\right)$.
-In this article, we study the large deviation probabilities for the maximum of the two-type reducible BBM. That is the convergence rate of the following probability
+In this article, we study the large deviation probabilities for the maximum of a two-type reducible BBM. That is the convergence rate of the following probability
 $\mathbb{P}(M_{t}>\theta m_{t}),\quad\theta>1,$
@@
 Theorem 1.1.
-If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$, then the following limit exists for any $\theta>1$,
+Assume $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$. Then the following limit exists for any $\theta>1$,
 $\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}}\theta t\right)=A,$
-where for $\beta<1$,
+where if $\beta\leq 1$, or $\beta>1$ and $\sigma^{2}\geq 1$, then
 $A=\beta\left(1-\theta^{2}\right),$
-for $\beta>1$ and $\sigma^{2}\geq 1$, we have
-$A=\beta\left(1-\theta^{2}\right),$
-for $\beta>1$ and $\sigma^{2}<1$, we have
-$A=\begin{cases}\beta\left(1-\theta^{2}\right),&\theta^{2}\leq\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta},\\ 1-\theta^{2}\beta\sigma^{2},&\theta^{2}\geq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta},\\ \frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.\end{cases}$
+and if $\beta>1$ and $\sigma^{2}<1$, then we have
+$A=\begin{cases}\beta\left(1-\theta^{2}\right),&\theta^{2}\leq\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta},\\ 1-\theta^{2}\beta\sigma^{2},&\theta^{2}\geq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta},\\ \frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}<\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.\end{cases}$
 Theorem 1.2.
-If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$, then the following limits exists for any $\theta>1$,
+Assume $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$. Then the following limit exists for any $\theta>1$,
 $\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)=:A,$
-where for $\beta>1$, we have
+where if $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, we have
 $A=1-\theta^{2},$
-for $\beta\leq 1,\sigma^{2}\leq 1$, we have
-$A=1-\theta^{2},$
-for $\beta\leq 1,\sigma^{2}>1$, we have
+and if $\beta\leq 1,\sigma^{2}>1$, we have
 $A=\begin{cases}1-\theta^{2},&\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)},\cr\beta-\frac{\theta^{2}}{\sigma^{2},}&\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4},\cr\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}<\theta^{2}<\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.\end{cases}$
 Theorem 1.3.
-If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$, then for any $\theta>1$, we have
+Assume $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$. Then for any $\theta>1$, we have
 $\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\xi\theta t\right)=\begin{cases}1-\frac{\xi^{2}\theta^{2}}{2},&\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}},\\ \beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}},&\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}},\cr\frac{\sqrt{2}(1-\beta)-\theta^{2}\left(\sigma^{2}-\beta\right)^{2}-1}{\sqrt{2}\left(\sigma^{2}-1\right)}+1,&\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}},\end{cases}$
@@
 where $C(\theta,\bar{u})$ is a constant depending only on $\theta$ and $\bar{u}$; see [11], [2] and [5].
-In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of two-type inreduciable BBM.
-It is possible to obtain more preciser form as (3) with the help of truncated second moment method and the F-KPP equation.
+In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of a two-type inreduciable BBM.
+It is possible to obtain a form as (3) with the help of the so-called truncated second moment method and the F-KPP equation.
 Our method may also apply to study the cases of $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{II}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}}\cap\mathcal{C}_{\text{III}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{III}}$; see [17] for related results.
@@
 We have three cases.
-i).
+Case I.
 If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\leq 1$, then $u\mapsto f_{1}^{\prime}(u)$ is an increasing function.
@@
 (15) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
-ii). If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\geq 1,\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
+Case II. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\geq 1,\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
 (16) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
-iii). If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
+Case III. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
 First, if $\theta^{2}<\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}$, which implies $f^{\prime}(1)>0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and
@@
 $\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)\right\}.$
-First, from 3, we have
+First, from (3), we have
 $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
@@
 $f_{2}^{\prime}(u)=\beta-1+\frac{\theta^{2}}{\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}\left(\sigma^{2}-1\right).$
-Solving $f^{\prime}(u_{0})=0$, we have
+Solving $f_{2}^{\prime}(u_{0})=0$, we have
 (39) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\theta-1}{\left|\sigma^{2}-1\right|},$
@@
 $\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\right\}.$
-First, from 3, we have
+First, from (3), we have
 $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2}\theta t\right)$
@@
 4.2.1. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(0)$.
-This corresponds to the cases that $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, or $\beta\leq 1,\sigma^{2}>1$ and $\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}.$ The arguments are parallel to those in Section 3.2.1.
-We only give an outline.
-Choose
+This corresponds to the cases that $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, or $\beta\leq 1,\sigma^{2}>1$ and $\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}.$ Choose
 $I_{t}=[\varepsilon t,\varepsilon t+1],\quad x_{t}=-\varepsilon t,$
@@
 4.2.2. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(1)$.
-This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Again, we only give an outline of the proof.
-Choose
+This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Choose
 $I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2}\theta t+\varepsilon t,$
@@
 $f_{3}^{\prime}(u)=\beta-1+\frac{\xi^{2}\left(\sigma^{2}-1\right)\theta^{2}}{2\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}.$
-Solving $f^{\prime}(u_{0})=0$ yields
+Solving $f_{3}^{\prime}(u_{0})=0$ yields
 (51) $u_{0}=\frac{\xi\theta\sqrt{\frac{1-\sigma^{2}}{2(\beta-1)}}-1}{\sigma^{2}-1}$
@@
 $\mathbb{P}\left(M_{t}\geqslant\xi\theta t\right)\geqslant\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right),\,\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)\right\}.$
-First, from 3, we have
+First, from (3), we have
 $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\xi\theta t\right)$

### v4

@@
 We consider a two-type reducible branching Brownian motion, defined as a particle system on the real line in which particles of two types move according to independent Brownian motions and create offspring at a constant rate.
-Particles of type $1$ can give birth to particles of types $1$ and $2$, but particles of type $2$ only give birth to descendants of type $2$.
+Particles of type $1$ can give birth to particles of types $1$ and $2$, but particles of type $2$ only give births to descendants of type $2$.
 Under some specific conditions, Belloum and Mallein in [5] showed that the maximum position $M_{t}$ of all particles alive at time $t$, suitably centered by a deterministic function $m_{t}$, converge weakly.
@@
 $\mathbb{P}(M_{t}\geq\theta m_{t}),\quad\theta>1.$
-We shall show that the decay rate function exhibits phase transitions depending on certain relations between $\theta$, the variance of the underling Brownian motion and the branching rate.
+We shall show that the decay rate function exhibits phase transitions depending on certain relations between $\theta$, the variance of the underlying Brownian motion and the branching rate.
 Key words and phrases:
@@
 On the other hand, it is the fundamental model to understand the BBM-universality class which includes the 2-dim Gaussian free field [8, 9] and 2-dim cover times [12], etc.
-An one-dimensional standard binary branching Brownian motion is a continuous-time particle system on the real line which can be constructed as follows.
+A one-dimensional standard binary branching Brownian motion is a continuous-time particle system on the real line which can be constructed as follows.
 It starts with one individual located at the origin at time 00 that moves according to a Brownian motion with variance $\sigma^{2}>0$.
@@
 Up to a dilation of time and space, in this work we shall always assume that $\beta_{2}=\sigma_{2}=1$.
-This model have been studied extensively by many authors; see [7, 14, 15, 17].
-With abuse of notation, for all $t\geq 0$, we denote by ${\mathcal{N}}_{t}$ the collection of the individuals alive at time $t$ for both the standard BBM and the two-type reducible BBM.
+This model has been studied extensively by many authors; see [7, 14, 15, 17].
+With abuse of notation, for all $t\geq 0$, we denote by ${\mathcal{N}}_{t}$ the collection of individuals alive at time $t$ for both the standard BBM and the two-type reducible BBM.
 And denote by ${\mathcal{N}}_{t}^{1}$ and ${\mathcal{N}}_{t}^{2}$ the set of particles of type 1 and type 2 at time $t$, respectively.
@@
 $M_{t}:=\max\{X^{u}_{t}:u\in{\mathcal{N}}(t)\}.$
-To distinguish the two-type reducible BBM and the standard BBM, we shall regard the standard BBM as $(\sigma^{2},\beta)$-BBM to emphasis the variance of the Brownian motion and the branching rate.
+To distinguish the two-type reducible BBM and the standard BBM, we shall regard the standard BBM as $(\sigma^{2},\beta)$-BBM to emphasize the variance of the Brownian motion and the branching rate.
 We also denote by $\mathbb{P}_{\sigma^{2},\beta}$ its law and write $\bar{X}^{u}_{s}$ for the position at time $s$ of the individual $u$ of a $(\sigma^{2},\beta)$-BBM.
@@
 $\frac{1}{2}\bar{\omega}^{\prime\prime}+\sqrt{2}\bar{\omega}^{\prime}-\bar{\omega}(1-\bar{\omega})=0.$
-Lalley and Sellke [16] characterised $\bar{\omega}$ in terms of a random shift of the Gumbel distribution by introducing the so-called derivative martingale.
+Lalley and Sellke [16] characterized $\bar{\omega}$ in terms of a random shift of the Gumbel distribution by introducing the so-called derivative martingale.
 Parallel results of (2) for the two-type reducible BBM were obtained in [5].
-To state their results, we need to decompose the state space $(\beta,\sigma^{2})\in{\mathbb{R}}_{+}^{2}$ into three regions:
+To state their results, we need to decompose the state space $(\beta,\sigma^{2})\in{\mathbb{R}}_{+}^{2}$ into three regions (see Figure 1 below for the phase diagram):
 $\mathcal{C}_{\text{I}}$ =$=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}>\frac{\mathbb{1}_{\{\beta\leq 1\}}}{\beta}+\mathbb{1}_{\{\beta>1\}}\frac{\beta}{2\beta-1}\right\},$
@@
 $\mathcal{C}_{\text{III}}$ =$=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}+\beta>2\text{ and }\sigma^{2}<\frac{\beta}{2\beta-1}\right\}.$
+Figure 1. Phase diagram of the two-type reducible BBM
 From Belloum and Mallein [5], we have:
@@
 Finally, if $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$, the speed of the process is a mixture of the long-time asymptotic of the processes of particles of type $1$ and $2$, which is larger than $\max\left(\sqrt{2},\sqrt{2\beta\sigma^{2}}\right)$.
+We remark here that before the work of Belloum and Mallein [5], Biggins’ work [7] was the first to consider the maximum of multi-type reducible branching random walks.
+After the work of Belloum and Mallein [5], as $\mathcal{C}_{I},\mathcal{C}_{II},\mathcal{C}_{III}$ does not cover all the possible values for $(\beta,\sigma^{2})\in(0,\infty)^{2}$, Belloum [4] studied the case of $\beta=\sigma^{2}=1$.
+Later, Ma and Ren [17] studied the other cases and completed the phase diagram.
 In this article, we study the large deviation probabilities for the maximum of a two-type reducible BBM. That is the convergence rate of the following probability
@@
 Assume $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$. Then the following limit exists for any $\theta>1$,
-$\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}}\theta t\right)=A,$
-where if $\beta\leq 1$, or $\beta>1$ and $\sigma^{2}\geq 1$, then
+$\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)=A,$
+where if $\sigma^{2}\geq 1$, then
 $A=\beta\left(1-\theta^{2}\right),$
@@
 $\lim_{t\to\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)=:A,$
-where if $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, we have
+where if $\sigma^{2}\leq 1$, we have
 $A=1-\theta^{2},$
 and if $\beta\leq 1,\,\sigma^{2}>1$, we have
-$A=\begin{cases}1-\theta^{2},&\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)},\cr\beta-\frac{\theta^{2}}{\sigma^{2},}&\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4},\cr\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}<\theta^{2}<\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.\end{cases}$
+$A=\begin{cases}1-\theta^{2},&\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)},\cr\beta-\frac{\theta^{2}}{\sigma^{2},}&\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4},\cr\frac{2\sqrt{(1-\beta)\left(\sigma^{2}-1\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}<\theta^{2}<\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.\end{cases}$
 Theorem 1.3.
@@
 where $C(\theta,\bar{u})$ is a constant depending only on $\theta$ and $\bar{u}$; see [11], [2] and [6].
-In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of a two-type inreduciable BBM.
+In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of a two-type reducible BBM.
 It is possible to obtain a form as (3) with the help of the so-called truncated second moment method and the F-KPP equation.
-Our method may also apply to study the cases of $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{II}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}}\cap\mathcal{C}_{\text{III}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{III}}$; see [17] for related results.
-In the next section, we will present some technical lemmas borrowed from existed literatures. The proofs for Theorems 1.1, 1.2 and 1.3 will be given in Sections 3, 4 and 5, respectively.
+Our method may also apply to study the cases of $(\beta,\sigma^{2})\in\partial\mathcal{C}_{\text{I}}\cap\partial\mathcal{C}_{\text{II}}$ or $(\beta,\sigma^{2})\in\partial\mathcal{C}_{\text{II}}\cap\partial\mathcal{C}_{\text{III}}$ or $(\beta,\sigma^{2})\in\partial\mathcal{C}_{\text{I}}\cap\partial\mathcal{C}_{\text{III}}$; see [17] for related results.
+In the next section, we will present some technical lemmas borrowed from the existing literature. The proofs for Theorems 1.1, 1.2 and 1.3 will be given in Sections 3, 4 and 5, respectively.
 2. Technical lemmas
 Lemma 2.1.
-([13], Lemma 2.1) Let $X$ be a centered Gaussian random variables with variance $\sigma^{2}>0$. Then
+([13], Lemma 2.1) Let $X$ be a centered Gaussian random variable with variance $\sigma^{2}>0$. Then
 (5) $\mathbb{P}(X>a)=(2\pi)^{-1/2}(\sigma/a)e^{-\frac{a^{2}}{2\sigma^{2}}}\left(1+O(\sigma^{2}/a^{2})\right),\quad\text{as }a/\sigma\rightarrow\infty,$
 with the r.h.s. above without error term being an upper bound valid for any $a>0.$
-Furthermore, for $(\sigma^{2},\beta)$-BBM and $10<x<$ 1, $\mathbb{P}_{\sigma^{2},\beta}-a.s.$,
-(6) $\frac{1}{t}\log{\mathcal{N}}_{t}\left(\sqrt{2\beta}x\sigma t)\right)\overset{t\rightarrow\infty}{\longrightarrow}(1-x^{2})\beta,$
-where ${\mathcal{N}}_{t}(x)$ denote the number of particles, in the $(\sigma^{2},\beta)$-BBM, alive at time $t$ and positioned in $[tx,\infty)$; see [6].
 If $u\in\mathcal{N}_{t}^{2}$, we denote by $T(u)$ the time at which the oldest ancestor of type 2 of $u$ was born.
-The set $\mathfrak{B}$ of particles of type $2$ that are born from a particle of type $1$, that can be defined as
+The set $\mathfrak{B}$ of particles of type $2$ that are born from a particle of type $1$, which can be defined as
 $\mathfrak{B}:=\{u\in\cup_{t\geq 0}{\mathcal{N}}^{2}_{t}:T(u)=b_{u}\},$
@@
 $\mathbb{E}\left[\sum_{u\in{\mathcal{N}}_{t}^{2}}g((X_{s}^{u},s\leq t),T(u))\right]$
-$\quad=\alpha\int_{0}^{t}e^{\beta s+(t-s)}E\left[g((\sigma B_{u\wedge s}+(B_{u}-B_{u\wedge s}),u\leq t),s)\right]{\rm d}s.$
+$\quad=\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{E}\left[g((\sigma B_{u\wedge s}+(B_{u}-B_{u\wedge s}),u\leq t),s)\right]{\rm d}s.$
 Proof.
-The first result is just the usual many-to-one lemma. The second one can be found in Proposition 4.1 in [4]. ∎
+The first result is just the usual many-to-one lemma. The second can be found in Proposition 4.1 in [5]. ∎
 3. Proof of Theorem 1.1
@@
 $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}>\frac{\mathbb{1}_{\{\beta\leq 1\}}}{\beta}+\frac{\beta\mathbb{1}_{\{\beta>1\}}}{2\beta-1}\right\}.$
+Before going into the details, we would like to give a rough explanation about the general strategy to all proofs.
+Our method is quite simple and direct.
+For the upper bound, we just use the first moment argument.
+Then the task is reduced to understand the birth time $T(u)$ of the particle whose position at time $t$ is larger than $\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t.$ For the lower bound, we may think that oldest ancestor of type $2$ of the maximal particle at time $t$ was born at some time $u_{0}t$ in a position around $x_{0}t$ for some $u_{0}$ and $x_{0}$.
+So the limit $A$ is consist of two parts: the effort of some type $1$ particles to achieve position $x_{0}t$ at time $u_{0}t$ and the effort that some type $2$ particles to achieve position $\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{0}t$ at time $t-u_{0}t$.
+The same strategy also apply well to the other two cases.
 3.1. Upper bound
 Proof.
-From Lemma 2.3, we have
-(7) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
-(8) $=$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
-(9) $\leqslant$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)+\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2},X_{t}^{u}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)$
-(10) $\leqslant$ $2\max\left[e^{\beta t}\mathbb{P}\left(\sigma B_{t}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right),\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{P}\left(\sigma B_{s}+B_{t}-B_{s}\geqslant\sqrt{2\sigma^{2}\beta}\theta t\right)\mathrm{d}s\right]$
-(11) =:absent:=:= : $2\max\{I_{1},I_{2}\}.$
-Then by (5)
-$I_{1}$ $\leq O(1)\cdot\frac{1}{\sqrt{t}}\cdot e^{\beta t}\cdot e^{-\frac{2\beta\theta^{2}t^{2}}{2t}}$
+First,
+(6) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+(7) $=$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t},X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+(8) $\leqslant$ $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1},X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)+\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2},X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+[…]
+Then by Lemma 2.3 and (5), we have
+I $\leq e^{\beta t}\mathbb{P}\left(\sigma B_{t}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+$\leq O(1)\cdot\frac{1}{\sqrt{t}}\cdot e^{\beta t}\cdot e^{-\frac{2\beta\theta^{2}t^{2}}{2t}}$
 $=\frac{O(1)}{\sqrt{t}}\cdot e^{\beta\left(1-\theta^{2}\right)t}.$
-Noticing that
-$\operatorname{Var}\left(\sigma B_{s}+B_{t}-B_{s}\right)=\sigma^{2}\cdot s+(t-s),$
-we have
-(12) $I_{2}$ $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\sigma^{2}\beta\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}\mathrm{d}s$
-(13) $=\alpha\int_{0}^{1}e^{f(u)t}\cdot\frac{\sqrt{t}}{\sqrt{2\pi\left((\sigma^{2}-1)u+1\right)}}\mathrm{d}u,$
-where
-(14) $f_{1}(u)=(\beta-1)u+1-\frac{\sigma^{2}\beta\theta^{2}}{(\left.\sigma^{2}-1\right)u+1}.$
-Notice that $f_{1}(1)=\beta(1-\theta^{2})$. Then our task is reduced to know $\max_{0\leq u\leq 1}f(u)$. To end this, let us consider
-$f_{1}^{\prime}(u)=\beta-1+\frac{\sigma^{2}\beta\theta^{2}}{\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}\left(\sigma^{2}-1\right).$
-We have three cases.
-Case I.
-If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\leq 1$, then $u\mapsto f_{1}^{\prime}(u)$ is an increasing function.
-In this case, by noticing that $\beta\sigma^{2}>1$, one has
-$f_{1}^{\prime}(0)=\beta-1+\sigma^{2}\beta(\sigma^{2}-1)=\beta(1-\sigma^{2})+\sigma^{2}\beta\sigma^{2}-1>\beta(1-\sigma^{2})+\sigma^{2}-1\geq 0\cdot$
-Thus $f_{1}^{\prime}(u)\geq 0$ for all $u\in[0,1]$. Hence $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and then we have
-(15) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
-Case II. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\geq 1,\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
-(16) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
-Case III. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
-First, if $\theta^{2}<\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}$, which implies $f^{\prime}(1)>0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and
-(17) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\beta\left(1-\theta^{2}\right).$
-Second, if $\theta^{2}>\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$,  which means $f_{1}^{\prime}(0)<0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(0)$ and
-(18) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq 1-\theta^{2}\beta\sigma^{2}.$
-Last, if $\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$, then $f_{1}^{\prime}(0)\geq 0$ and $f^{\prime}(1)\leq 0$.
-Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$, where $f_{1}^{\prime}(u_{0})=0$.
-And one can check that
-(19) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\sigma\theta\cdot\sqrt{\beta}-1}{\left|\sigma^{2}-1\right|},$
-and
-$f_{1}(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
-Therefore,
-(20) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2\sigma^{2}\beta}\theta t\right)\leq\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
-This proves the upper bound. ∎
-3.2. Lower bound for Theorem 1.1
-Notice that
-$\mathbb{P}\left(M_{t}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-$\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)\right\}.$
-First, from (3), we have
-$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-$=\beta\left(1-\theta^{2}\right).$
 On the other hand, set
@@
 According to construction of our model, we have
-(21) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-(22) $=\mathbb{P}\left(\exists u\,\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-(23) $=1-\mathbb{P}\left(\forall\,u\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}<\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-(24) $=1-\mathbb{E}\left[\prod_{u\in\mathfrak{B}_{t}}\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right]$
-(25) $=1-\mathbb{E}\left[\exp\left(-\sum_{u\in\mathfrak{B}_{t}}-\ln\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right)\right],$
+$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+$=\mathbb{P}\left(\exists u\,\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+$=1-\mathbb{P}\left(\forall\,u\in\mathfrak{B}_{t}:X_{T(u)}^{u}+M_{t-T(u)}^{u}<\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+$=1-\mathbb{E}\left[\prod_{u\in\mathfrak{B}_{t}}\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right]$
+$=1-\mathbb{E}\left[\exp\left(-\sum_{u\in\mathfrak{B}_{t}}-\ln\varphi_{t}\left(X_{T(u)}^{u},T(u)\right)\right)\right],$
 where
-$1-\varphi_{t}(x,s)=\bar{u}\left(t-s,\,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x\right)=\mathbb{P}_{1,1}\left(\bar{M}_{t-s}\geq\sqrt{2\sigma^{2}\theta^{2}\beta}\cdot t-x\right).$
+$1-\varphi_{t}(x,s)=\bar{u}\left(t-s,\,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x\right)=\mathbb{P}_{1,1}\left(\bar{M}_{t-s}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x\right).$
 and $\left\{M^{u}_{t-T(u)}:u\in\mathfrak{B}_{t}\right\}$ are independent random variables such that $M^{u}_{t-s}$ has the same law as $\bar{M}_{t-s}$ under $\mathbb{P}_{1,1}$.
-Applying lemma 2.2, we have
-(26) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-(27) $=1-\mathbb{E}\left[\exp\left\{-\alpha\int_{0}^{t}\sum_{u\in{\mathcal{N}}_{s}^{1}}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right\}\right]$
-(28) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right\}\right]$
-(29) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right]$
-(30) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right],$
-where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$, and
+Applying Lemmas 2.2, we have
+(10) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+(11) $=1-\mathbb{E}\left[\exp\left\{-\alpha\int_{0}^{t}\sum_{u\in{\mathcal{N}}_{s}^{1}}\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right\}\right].$
+With Lemma 2.3 in hand, applying elementary inequality $1-e^{-x}\leq x$ for $x\geq 0$, we have
+(12) II $\leq\mathbb{E}\left[\alpha\int_{0}^{t}\sum_{u\in{\mathcal{N}}_{s}^{1}}\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-X_{s}^{u}\right){\rm d}s\right]$
+(13) $=\alpha\int_{0}^{t}{\rm d}s\int_{-\infty}^{+\infty}e^{\beta s}\,\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x\right)\frac{1}{\sqrt{2\pi\sigma^{2}s}}e^{-\frac{x^{2}}{2\sigma^{2}s}}{\rm d}x$
+(14) $=\alpha t^{3/2}\int_{0}^{1}{\rm d}s\int_{-\infty}^{+\infty}e^{\beta st}\,\bar{u}\left((1-s)t,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-xt\right)\frac{1}{\sqrt{2\pi\sigma^{2}s}}e^{-\frac{x^{2}t}{2\sigma^{2}s}}{\rm d}x.$
+By (3), for any $10<s<$ 1 and $x<\sqrt{2\beta\sigma^{2}\theta^{2}}-\sqrt{2}(1-s)$, we have, as $t\rightarrow\infty$,
+$\bar{u}\left((1-s)t,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-xt\right)\leq O(1)\frac{1}{\sqrt{t}}\exp{\left\{\left(1-s-\frac{(\sqrt{2\beta\sigma^{2}\theta^{2}}-x)^{2}}{2(1-s)}\right)t\right\}}.$
+Thus
+$I_{2}\leq O(1)t\int_{0}^{1}{\rm d}u\int_{-\infty}^{+\infty}e^{f_{1}^{*}(u,x)t}\frac{1}{\sqrt{u}}{\rm d}x,$
+where for $0<u<1,\,x\in{\mathbb{R}},$
+$f^{*}_{1}(u,x)=\beta u-\frac{x^{2}}{2\sigma^{2}u}+\left(1-u-\frac{(\sqrt{2\beta\sigma^{2}\theta^{2}}-x)^{2}}{2(1-u)}\right){\bf 1}_{\{x<\sqrt{2\beta\sigma^{2}\theta^{2}}-\sqrt{2}(1-u)\}},$
+By studying function $x\mapsto f^{*}_{1}(u,x)$, one can see that
+(15) $\max_{x\in{\mathbb{R}}}f^{*}_{1}(u,x)$ $=$ $f^{*}_{1}(u,x^{*}(u))=:f_{1}(u)$
+(16) $=$ $\begin{cases}(\beta-1)u+1-\frac{\beta\sigma^{2}\theta^{2}}{1-u+\sigma^{2}u},\quad&\text{if }u(\sigma^{2}-1)<\sqrt{\beta\sigma^{2}\theta^{2}}-1,\cr\beta u-\frac{(\sqrt{\beta\sigma^{2}\theta^{2}}-(1-u))^{2}}{\sigma^{2}u},\quad&\text{if }u(\sigma^{2}-1)\geq\sqrt{\beta\sigma^{2}\theta^{2}}-1,\end{cases}$
+with
+$x^{*}(u)=\begin{cases}\frac{\sqrt{2\beta}\theta\sigma^{3}u}{1-u+\sigma^{2}u},\quad&\text{if }u(\sigma^{2}-1)<\sqrt{\beta\sigma^{2}\theta^{2}}-1,\cr\sqrt{2\beta\sigma^{2}\theta^{2}}-\sqrt{2}(1-u),\quad&\text{if }u(\sigma^{2}-1)\geq\sqrt{\beta\sigma^{2}\theta^{2}}-1,\end{cases}$
+where $u(\sigma^{2}-1)<\sqrt{\beta\sigma^{2}\theta^{2}}-1$ is equivalent to $\frac{\sqrt{2\beta}\theta\sigma^{3}u}{1-u+\sigma^{2}u}<\sqrt{2\beta\sigma^{2}\theta^{2}}-\sqrt{2}(1-u).$ Furthermore,
+$f^{\prime}_{1}(u)=\begin{cases}\beta-1+\frac{\beta\sigma^{2}\theta^{2}(\sigma^{2}-1)}{(1-u+\sigma^{2}u)^{2}},\quad&\text{if }u(\sigma^{2}-1)<\sqrt{\beta\sigma^{2}\theta^{2}}-1,\cr\frac{(\beta\sigma^{2}-1)u^{2}+(\sqrt{\beta\sigma^{2}\theta^{2}}-1)^{2}}{\sigma^{2}u^{2}},\quad&\text{if }u(\sigma^{2}-1)\geq\sqrt{\beta\sigma^{2}\theta^{2}}-1.\end{cases}$
+Also notice that for $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$, we always have
+$\beta\sigma^{2}\geq\begin{cases}1,\quad&\beta\leq 1,\cr\frac{\beta^{2}}{2\beta-1}\geq 1,\quad&\beta>1.\end{cases}$
+Then we have two cases.
+Case I. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$ for $10<u<$ 1. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
+$\max_{0\leq u\leq 1,\,x\in\mathbb{R}}f^{*}_{1}(u,x)=f_{1}\left(1\right)=\beta(1-\theta^{2}).$
+Hence
+(17) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq\beta\left(1-\theta^{2}\right).$
+Case II. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
+First, if $\theta^{2}<\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}$, which implies $f_{1}^{\prime}(1)>0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and
+(18) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq\beta\left(1-\theta^{2}\right).$
+Second, if $\theta^{2}>\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$,  which means $f_{1}^{\prime}(0)<0$, then $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(0)$ and
+(19) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq 1-\theta^{2}\beta\sigma^{2}.$
+Last, if $\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}$, then one can check that $f_{1}^{\prime}(0)\geq 0$ and $f_{1}^{\prime}(1)\leq 0$.
+Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$ with
+(20) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\sigma\theta\cdot\sqrt{\beta}-1}{\sigma^{2}-1}.$
+Therefore,
+(21) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq f_{1}(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
+This proves the upper bound.
+Remark. Notice that
+$\operatorname{Var}\left(\sigma B_{s}+B_{t}-B_{s}\right)=\sigma^{2}\cdot s+(t-s).$
+Then applying Lemma 2.3 and (5), we also have
+(22) $I_{2}$ $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{P}\left(\sigma B_{s}+B_{t}-B_{s}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\mathrm{d}s$
+(23) $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\beta\sigma^{2}\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}\mathrm{d}s$
+(24) $=\alpha\int_{0}^{1}e^{f_{1}(u)t}\cdot\frac{\sqrt{t}}{\sqrt{2\pi\left((\sigma^{2}-1)u+1\right)}}\mathrm{d}u.$
+If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$, $\beta\leq 1$ and $\theta^{2}\leq\frac{(1-\beta)\sigma^{2}}{(\sigma^{2}-1)\beta}$, then one can check that $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$.
+So we will get
+$\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\leq f^{\prime}(u_{0}),$
+which is larger than $\beta(1-\theta^{2}).$ Thus we would get a rough upper bound. ∎
+3.2. Lower bound for Theorem 1.1
+Notice that
+(25) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+(26) $\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)\right\}.$
+First, from (3), we have
+(27) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+(28) $=\beta\left(1-\theta^{2}\right).$
+Recall (10). Then
+(29) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+(30) $\geq 1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right]$
+(31) $=1-\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right],$
+where $I_{t}\subset[0,t]$ is an interval of length $1$, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$, and
 ${\mathcal{N}}_{s}^{1}(a)=\{u\in{\mathcal{N}}_{s}^{1}:X_{s}^{u}\geq a\}.$
-Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Furthermore, by Jensen’s inequality and Fubini’s theorem, we have
-(31) $\mathbb{E}\left[\exp\left\{-\alpha\int_{I_{t}}\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right){\rm d}s\right\}\right]$
-(32) $\leq\int_{I_{t}}\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s.$
+Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Then, applying Jensen’s inequality and Fubini’s theorem, gives
+(32) $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+(33) $\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s.$
 3.2.1. The case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(0)$.
@@
 for $\varepsilon>0$ small enough. Then for $s\in I_{t}$,
-$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
-$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(-\varepsilon t)}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\},|{\mathcal{N}}_{s}^{1}(-\varepsilon t)|\geq 1\right]$
+$\mathbb{E}\left[\exp\left\{-\alpha\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right)\right\}\right]$
+$\leq\mathbb{E}\left[\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\right\},|{\mathcal{N}}_{s}^{1}(-\varepsilon t)|\geq 1\right]$
 $\quad+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(-\varepsilon t)|=0\right)$
-$\leq\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right)+\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}<-\varepsilon t\right).$
-Therefore,
-$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-$\geq 1-\int_{I_{t}}\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s$
-$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right){\rm d}s$
-$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-\varepsilon t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right){\rm d}s.$
+$=\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\right\}\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right)+\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}<-\varepsilon t\right).$
+This, together with (32), gives
+$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right){\rm d}s$
+$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-\varepsilon t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\right\}\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right){\rm d}s.$
 Then applying (3), we have
-$\lim_{\varepsilon\rightarrow 0}\lim_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t+\varepsilon t\right)\geq 1-\sigma^{2}\beta\theta^{2},$
+$\lim_{\varepsilon\rightarrow 0}\lim_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-\varepsilon t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t+\varepsilon t\right)\geq 1-\beta\sigma^{2}\theta^{2},$
 uniformly for $s\in[0,1]$. We further notice that
 $\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}\geq-\varepsilon t\right)\rightarrow 1,\quad\text{as }t\rightarrow\infty,$
-uniformly for $s\in[0,1]$. Then applying Jensen’s inequality and Fatou’s lemma, one has
-$\liminf_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$ $=1-\sigma^{2}\beta\theta^{2}$
-$\geq\beta(1-\theta^{2}),$
+uniformly for $s\in[0,1]$. Then applying Jensen’s inequality and Fatou’s lemma again, one has
+$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)=1-\beta\sigma^{2}\theta^{2}\geq\beta(1-\theta^{2}),$
 where the last inequality holds if $\theta^{2}\geq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\beta}$.
@@
 $\quad\bigcup\left\{\left(\theta^{2},\beta,\sigma^{2}\right):\theta^{2}\leq\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta},\,\beta>1,\,\frac{\beta}{2\beta-1}<\sigma^{2}<1\right\}.$
+One can make use (25) and (27) directly to conclude. We are done.∎
+3.2.3. The case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$.
+This corresponds to the case that $\beta>1,\sigma^{2}<1$, and
+$\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.$
+Recall $u_{0}$ from (20). We also recall that
+$f_{1}(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}\geq f_{1}(1)=\beta(1-\theta^{2}).$
 Choose
-$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2\beta\sigma^{2}}\theta t+\varepsilon t,$
-for $\varepsilon>0$ small enough. Then for $s\in I_{t}$,
-$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
-$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,-\varepsilon t\right)\right\},|{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right]$
-$\quad+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|=0\right).$
-Notice that
-$\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)=\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right).$
+$I_{t}=[u_{0}t,u_{0}t+1],\quad x_{t}=x_{0}t,$
+for some $0<x_{0}<\sqrt{2\beta\sigma^{2}\theta^{2}}$, whose value will be determined later. Then for $\varepsilon>0$ small enough and $s\in I_{t}$,
+$\mathbb{E}\left[\exp\left\{-\alpha\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{t}\right)\right\}\right]$
+$\leq\exp\left\{-\alpha\bar{u}\left(t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{0}t\right)\right\}\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{0}t)|\geq 1\right)+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{0}t)|=0\right),$
+where
+$x_{0}=\frac{\sqrt{2\beta}\theta u_{0}\sigma^{3}}{1-u_{0}+u_{0}\sigma^{2}}.$
 Therefore,
-$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right){\rm d}s$
-$\geq\int_{I_{t}}\left(1-\exp\left\{-\alpha\bar{u}\left(t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{s}>x_{t}\right){\rm d}s$
-$=\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)\right\}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>x_{t}\right){\rm d}s.$
-Since
-$\lim_{t\rightarrow\infty}\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)=1$
+$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{0}t\right)\right\}\right)$
+×$\qquad\quad\times\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{0}t)|\geq 1\right){\rm d}s.$
+Furthermore, the fact that $f_{1}^{\prime}(u_{0})=0$ gives
+$1-u_{0}+u_{0}\sigma^{2}=\sqrt{\frac{1-\sigma^{2}}{\beta-1}}\sqrt{\beta}\sigma\theta.$
+Then
+$x_{0}=\sqrt{2}\sigma^{2}\sqrt{\frac{\beta-1}{1-\sigma^{2}}}u_{0}>\sqrt{2\beta}\sigma u_{0}$
 and
-$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>x_{t}\right)=\beta(1-\theta^{2}),$
-uniformly for $s\in[0,1]$, then we have
-$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)=\beta(1-\theta^{2}).$
-3.2.3. The case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$.
-Notice that where $f_{1}^{\prime}(u_{0})=0$. This corresponds to the case that $\beta>1,\sigma^{2}<1$, and
-$\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.$
-We also recall that
-$f(u_{0})=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
-For $\varepsilon>0$ small enough, choose
-$I_{t}=[u_{0}t,u_{0}t+1],\quad x_{t}=x_{0}t,$
-for some $0<x_{0}<\sqrt{2\beta\theta^{2}\sigma^{2}}$, whose value will be determined later. Then for $s\in I_{t}$,
-$\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]$
-$\leq\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\},|{\mathcal{N}}_{s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right]$
-$\quad+\mathbb{P}_{\sigma^{2},\beta}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|<e^{(\beta_{0}-\varepsilon)t}\right)$
-$\leq\exp\left\{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\}\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)$
-$\quad+\mathbb{P}\left(|{\mathcal{N}}_{s}^{1}(x_{t})|<e^{(\beta_{0}-\varepsilon)t}\right),$
-where
-$\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0,\quad\text{and}\quad x_{0}=\frac{\sqrt{2\beta}\theta u_{0}\sigma^{3}}{1-u_{0}+u_{0}\sigma^{2}}.$
+$\sqrt{2\beta\sigma^{2}\theta^{2}}-x_{0}=\sqrt{2\beta}\sigma\theta\frac{1-u_{0}}{1-u_{0}+u_{0}\sigma^{2}}=\sqrt{2}\sqrt{\frac{\beta-1}{1-\sigma^{2}}}(1-u_{0})>\sqrt{2}(1-u_{0}),$
+because of
+$\sigma^{2}\geq\frac{\beta}{2\beta-1}>2-\beta,\quad\text{for }\beta>1.$
 Therefore,
-$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-$\geq\int_{0}^{1}\left(1-\exp\left\{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)\right\}\right)$
-×$\qquad\quad\times\mathbb{P}_{\sigma^{2},\beta}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
-We also notice that
-$\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}>1,$
-since
-$\sigma^{2}\geq\frac{\beta}{2\beta-1}>2-\beta,$
-for $\beta>1$. Then applying (3) again, we have
-$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{0}t\right)$
-$\quad\geq\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
-uniform for $s\in[0,1]$. Furthermore, according to (6),
-$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
-Thus
-$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1.$
-Since $\varepsilon>0$ is arbitrary, then applying Jensen’s inequality and Fatou’s lemma, one has
-(33) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t\right)$
-(34) $\geq\beta_{0}+\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
+$\frac{\left(\sqrt{2\beta\sigma^{2}\theta^{2}}-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}>1,\quad\text{for }\beta>1.$
+Thus applying (3) again, we have
+$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t-x_{0}t\right)$
+$\quad\geq\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}\theta^{2}}-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
+uniform for $s\in[0,1]$ and
+$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{0}t)|\geq 1\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{u_{0}t+s}>x_{0}t\right)$
+$=\left(1-\frac{x_{0}^{2}}{{2\beta}\sigma^{2}u_{0}^{2}}\right)\beta u_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}},$
+uniform for $s\in[0,1]$. Then applying Jensen’s inequality and Fatou’s lemma, one has
+(34) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2\beta\sigma^{2}\theta^{2}}\cdot t\right)$
+(35) $\geq\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}+\left(1-\frac{\left(\sqrt{2\beta\sigma^{2}\theta^{2}}-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
 (36) $=(\beta-1)u_{0}+1-\frac{\sigma^{2}\beta\theta^{2}}{(\left.\sigma^{2}-1\right)u_{0}+1}$
-(36) $=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}$
-(37) $\geq\beta(1-\theta^{2})=f_{1}(1),$
-where the last inequality holds for the case that $\beta>1,\sigma^{2}<1$, and
-$\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.$
-This proves Theorem 1.1 for the above case.
+(37) $=\frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
+This proves Theorem 1.1 for the case of $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(u_{0})$. We are done.
 4. Proof of Theorem 1.2
@@
 $(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}}=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}<\frac{\mathbb{1}_{\{\beta\leq 1\}}}{\beta}+\mathbb{1}_{\{\beta>1\}}(2-\beta)\right\}.$
-The arguments are exactly the same to those in Section 3. One just to replace $\sqrt{2\beta\sigma^{2}}\theta$ by $\sqrt{2}\theta$. We only give an outline here.
+One can use arguments similar to those in Section 3. We just replace $\sqrt{2\beta\sigma^{2}}\theta$ by $\sqrt{2}\theta$. However, for the upper bound we have a simpler method.
 4.1. Upper bound for Theorem 1.2
@@
 $\text{I}\leq O(1)\cdot\frac{1}{\sqrt{t}}\cdot e^{\beta t}\cdot e^{-\frac{2\theta^{2}t^{2}}{2\sigma^{2}t}}=\frac{O(1)}{\sqrt{t}}\cdot e^{\left(\beta-\frac{\theta^{2}}{\sigma^{2}}\right)t}.$
-and
-$\text{II}\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}.$
-Set
+and by Lemma 2.3, we also have
+(38) II $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{P}\left(\sigma B_{s}+B_{t}-B_{s}\geqslant\sqrt{2\theta^{2}}\cdot t\right)\mathrm{d}s$
+(39) $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}\mathrm{d}s$
+(40) $=\alpha\int_{0}^{1}e^{f_{2}(u)t}\cdot\frac{\sqrt{t}}{\sqrt{2\pi\left((\sigma^{2}-1)u+1\right)}}\mathrm{d}u,$
+where
 (41) $f_{2}(u)=(\beta-1)u+1-\frac{\theta^{2}}{(\left.\sigma^{2}-1\right)u+1},\quad u\in(0,1).$
-Then
-$f_{2}^{\prime}(u)=\beta-1+\frac{\theta^{2}}{\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}\left(\sigma^{2}-1\right).$
-Solving $f_{2}^{\prime}(u_{0})=0$, we have
-(39) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\theta-1}{\left|\sigma^{2}-1\right|},$
-and
-$f_{2}(u_{0})=\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
-Then we have several subcases.
+Then we have several cases.
 Case I. $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$ and $\beta>1$.
 In this case $u\mapsto f_{2}^{\prime}(u)$ is increasing and $f_{2}^{\prime}(1)<0$. Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(0)$ and hence
-(40) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
+(42) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
 Case II. $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$ and $\beta\leq 1,\sigma^{2}<1$.
 In this case $f_{2}^{\prime}(u)<0$ for all $u\in[0,1]$, Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(0)$ and hence
-(41) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
+(43) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
 Case III. $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$ and $\beta\leq 1,\sigma^{2}\geq 1$. In this case, $u\mapsto f_{2}^{\prime}(u)$ is decreasing.
 When $\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}$, then $f_{2}^{\prime}(0)<0$. Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(0)$ and hence
-(42) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
+(44) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
 When $\theta^{2}>\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}$, then $f_{2}^{\prime}(1)>0$, Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(1)$ and hence
-(43) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq\beta-\frac{\theta^{2}}{\sigma^{2}}.$
+(45) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq\beta-\frac{\theta^{2}}{\sigma^{2}}.$
 When $\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}\leq\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}$, then $f_{2}^{\prime}(0)\geq 0$ and $f_{2}^{\prime}(1)\leq 0$.
-Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(u_{0})$ and hence
-(44) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
+Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(u_{0})$ with $f_{2}^{\prime}(u_{0})=0$.
+Then we have
+(46) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\theta-1}{\sigma^{2}-1}$
+and hence
+(47) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq f_{2}(u_{0})=\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
 4.2. Lower bound for Theorem 1.2
 Notice that
-$\mathbb{P}\left(M_{t}\geqslant\sqrt{2}\theta t\right)$
-$\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\right\}.$
+(48) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2}\theta t\right)$
+(49) $\geqslant\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\right\}.$
 First, from (3), we have
-$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2}\theta t\right)$
-$=\beta-\frac{\theta^{2}}{\sigma^{2}}.$
-By similar reasoning as (21), (26) and (31), we have
+(50) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2}\theta t\right)$
+(51) $=\beta-\frac{\theta^{2}}{\sigma^{2}}.$
+By similar reasoning as (29) and (32), we have
 $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
-$\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s,$
-where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$. Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Recall from (38) that
+$\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2}\theta t-x_{t}\right)\right\}\right]{\rm d}s,$
+where $I_{t}\subset[0,t]$ is an interval of length $1$, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$.
+Both $I_{t}$ and $x_{t}$ will be determined later from case to case.
+Recall from (41) that
 $f_{2}(u)=(\beta-1)u+1-\frac{\theta^{2}}{(\left.\sigma^{2}-1\right)u+1},\quad u\in(0,1).$
@@
 4.2.2. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(1)$.
-This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Choose
-$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2}\theta t+\varepsilon t,$
-for $\varepsilon>0$ small enough. Then we have
-$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
-$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\sqrt{2}\theta t+\varepsilon t\right){\rm d}s.$
-Since
-$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(M_{(1-\varepsilon)t+s}>\sqrt{2}\theta t+\varepsilon t\right)\geq\beta-\frac{\theta^{2}}{\sigma^{2}},$
-then we have
-$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\geq\beta-\frac{\theta^{2}}{\sigma^{2}}.$
-This proves Theorem 1.2 for the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$
+This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Then with (48) and (50) in hand, Theorem 1.2 for the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}$ follows readily.
 4.2.3. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(u_{0})$ for $10<u_{0}<$ 1.
@@
 Choose
-$x_{0}=\frac{\sqrt{2}\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t,\quad\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0.$
+$x_{0}=\frac{\sqrt{2}\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t.$
 One can check that $x_{0}<\sqrt{2}\theta$ since $u_{0}>0.$ Then,
 $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
-$\geq\int_{0}^{1}\left(1-e^{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\sqrt{2}\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
+$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(t-u_{0}t-s,\sqrt{2}\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq 1\right){\rm d}s.$
 We also notice that
@@
 $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\sqrt{2}\theta t-x_{0}t\right)\geq\left(1-\frac{\left(\sqrt{2}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
-uniform for $s\in[0,1]$. Furthermore, according to (6),
-$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
-Thus
-$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1,$
-uniform for $s\in(0,1)$. Then one has
+uniform for $s\in[0,1]$ and
+$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq 1\right)=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}},$
+uniform for $s\in[0,1]$. Then one has
 (52) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$
-(46) $\geq\beta_{0}+\left(1-\frac{\left(\sqrt{2}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
+(53) $\geq\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}+\left(1-\frac{\left(\sqrt{2}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
 (54) $=(\beta-1)u_{0}+1-\frac{\theta^{2}}{\left(\sigma^{2}-1\right)u_{0}+1}$
-(48) $=\frac{-2\sqrt{(1-\beta)\left(\sigma^{2}-1\right)}\theta-\beta+\sigma^{2}}{\sigma^{2}-1}$
-(49) $\geq\beta-\frac{\theta^{2}}{\sigma^{2}}=f_{2}(1),$
-where the last inequality holds for the case that $\beta<1,\sigma^{2}>1$, and
-$\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\leq\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$
+(55) $=\frac{-2\sqrt{(1-\beta)\left(\sigma^{2}-1\right)}\theta-\beta+\sigma^{2}}{\sigma^{2}-1}.$
 This proves Theorem 1.2 for the above case.
@@
 $(\beta,\sigma^{2})\in\mathcal{C}_{\text{III}}=\left\{\left(\beta,\sigma^{2}\right):\sigma^{2}+\beta>2\text{ and }\sigma^{2}<\frac{\beta}{2\beta-1}\right\}.$
-Recall $\xi=\frac{\sigma^{2}-\beta}{\sqrt{2(1-\sigma^{2})(\beta-1)}}.$ The arguments are also exactly the same to those in Section 4.
-One just to replace $\theta$ by $\frac{\xi\theta}{\sqrt{2}}$.
-Again, We only give an outline here.
+Recall $\xi=\frac{\sigma^{2}-\beta}{\sqrt{2(1-\sigma^{2})(\beta-1)}}.$ The arguments are also exactly the same as those in Section 4.
+One just replace $\theta$ by $\frac{\xi\theta}{\sqrt{2}}$.
+We give only an outline here.
 5.1. Upper bound
@@
 and
-$\text{II}\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{\xi^{2}\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}.\$
+$\text{II}\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{\xi^{2}\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}{\rm d}s.\$
 For II, set
@@
 $f_{3}^{\prime}(u)=\beta-1+\frac{\xi^{2}\left(\sigma^{2}-1\right)\theta^{2}}{2\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}.$
-Solving $f_{3}^{\prime}(u_{0})=0$ yields
-(51) $u_{0}=\frac{\xi\theta\sqrt{\frac{1-\sigma^{2}}{2(\beta-1)}}-1}{\sigma^{2}-1}$
 and
@@
 If $\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, then $f_{3}^{\prime}(0)\leq 0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(0)$ and hence
-(52) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\xi\theta t\right)\leq f_{3}(0)=1-\frac{\xi^{2}\theta^{2}}{2}.$
-If $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$, then $f^{\prime}(1)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(1)$ and hence
-(53) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\xi\theta t\right)\leq f_{3}(1)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
+(57) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\xi\theta t\right)\leq f_{3}(0)=1-\frac{\xi^{2}\theta^{2}}{2}.$
+If $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$, then $f_{3}^{\prime}(1)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(1)$ and hence
+(58) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\xi\theta t\right)\leq f_{3}(1)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
 If $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, then $f_{3}^{\prime}(0)>0$ and $f_{3}^{\prime}(1)<0$.
-Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ and hence
-(54) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln P\left(M_{t}>\xi\theta t\right)\leq f_{3}(u_{0})=\frac{\sqrt{2}(1-\beta)-\theta^{2}\left(\sigma^{2}-\beta\right)^{2}-1}{\sqrt{2}\left(\sigma^{2}-1\right)}+1.$
+Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ with $f_{3}^{\prime}(u_{0})=0$ yields that
+(59) $u_{0}=\frac{\xi\theta\sqrt{\frac{1-\sigma^{2}}{2(\beta-1)}}-1}{\sigma^{2}-1}$
+and
+(60) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\xi\theta t\right)\leq f_{3}(u_{0})=\frac{\sqrt{2}(1-\beta)-\theta^{2}\left(\sigma^{2}-\beta\right)^{2}-1}{\sqrt{2}\left(\sigma^{2}-1\right)}+1.$
 5.2. Lower bound for Theorem 1.3
 Notice that
-$\mathbb{P}\left(M_{t}\geqslant\xi\theta t\right)\geqslant\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right),\,\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)\right\}.$
+(61) $\mathbb{P}\left(M_{t}\geqslant\xi\theta t\right)\geqslant\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right),\,\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)\right\}.$
 First, from (3), we have
-$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\xi\theta t\right)$
-$=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
-By similar reasoning as (21), (26) and (31), we have
+(62) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\xi\theta t\right)$
+(63) $=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
+By similar reasoning as (29) and (32), we have
 $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$
 $\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\xi\theta t-x_{t}\right)\right\}\right]{\rm d}s,$
-where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$. Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Recall from (50) that
+where $I_{t}\subset[0,t]$ is an interval of length $1$, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$.
+Both $I_{t}$ and $x_{t}$ will be determined later from case to case.
+Recall from (56) that
 $f_{3}(u)=(\beta-1)u+1-\frac{\xi^{2}\theta^{2}}{2((\left.\sigma^{2}-1\right)u+1)},\quad u\in(0,1).$
@@
 We further notice that
-$\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right)\rightarrow 1,\quad\text{as }t\rightarrow\infty,$
+$\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right)\rightarrow 1,\quad\text{as }t\rightarrow\infty.$
 Then applying Jensen’s inequality and Fatou’s lemma, one has
 $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $\geq 1-\frac{\xi^{2}\theta^{2}}{2}.$
-This proves Theorem 1.3 for the cases that $\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, since $1-\frac{\xi^{2}\theta^{2}}{2}\geq\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}$ in this case.
-5.2.2. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(1)$.
-This case corresponds to the case that $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$. Choose
-$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2}\theta t+\varepsilon t,$
-for $\varepsilon>0$ small enough. Then we have
+This proves Theorem 1.3 for the cases $\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$.
+5.2.2. The case of $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(1)$.
+This case corresponds to the case $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$. Similarly as before, (61) and (62) yield Theorem 1.3 for this case.
+5.2.3. The case of $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ for $10<u_{0}<$ 1.
+Notice that $f_{3}^{\prime}(u_{0})=0$. This corresponds to the case $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$. Choose
+$x_{0}=\frac{\xi\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t.$
+Obviously $x_{0}<\xi\theta$, since $u_{0}>0.$ Then
 $\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$
-$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\xi\theta t+\varepsilon t\right){\rm d}s.$
-Since
-$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\xi\theta t+\varepsilon t\right)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}},$
-then we have
-$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
-This proves Theorem 1.3 for the case that $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$.
-5.2.3. The case of $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ for $10<u_{0}<$ 1.
-Notice that $f_{3}^{\prime}(u_{0})=0$. This corresponds to the case that $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$. Choose
-$x_{0}=\frac{\xi\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t,\quad\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0.$
-Obviously $x_{0}<\xi\theta$ since $u_{0}>0.$ Then
-$\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$
-$\geq\int_{0}^{1}\left(1-e^{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\,\xi\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
+$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(t-u_{0}t-s,\,\xi\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq 1\right){\rm d}s.$
 We also notice that
@@
 $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\bar{u}\left(t-u_{0}t-s,\xi\theta t-x_{0}t\right)\geq\left(1-\frac{\left(\xi\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right),$
-uniform for $s\in[0,1]$. Furthermore, according to (6),
-$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
-Thus
-$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1,$
-uniform for $s\in(0,1)$. Then one has
-(55) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=f_{3}(u_{0})$
-(56) $\geq\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}=f_{3}(1),$
-where the last inequality holds for the case that $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$.
-This proves Theorem 1.3 for the above case.
-We are done.
+uniform for $s\in[0,1]$ and
+$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq 1\right)=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}},$
+uniform for $s\in[0,1]$. Then one has
+(64) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)=f_{3}(u_{0}).$
+This proves Theorem 1.3 for the above case. We are done.
+Acknowledgement
+We are deeply indebted to Haojun Zhao who pointed out an error in the proof of Theorem 1.1 in the previous version.